Grade upto college level Physical Chemistry

A graph is plotted between PVm along Y-axis and P along X-axis, where Vm is the molar volume of a real gas. Find the intercept along Y-axis.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
The van der Waal equation (for one mole) of a real gas is
\left ( P + \frac{a}{V_{m}^{2}} \right )(Vm - b) = RT
PVm – Pb + a/V – ab/V2 = RT
PVm = RT + Pb – a/Vm + \frac{ab}{V_{m}^{2}} …..(i)
NOTE THIS STEP :To calculate the intercept P → 0, hence Vm → ∞ due to which the last two terms on the right side of the equation (i) can be neglected.
∴ PVm = RT + Pb
When P = 0, intercept = RT